Suppose a perfectly competitive industry can produce Roman candles at a constant marginal cost of R12 per unit. Once industry is monopolized, marginal cost rise to R16 per unit because R4 per unit must be paid to lobbyists to ensure that only this firm receives a Roman candle license. Suppose the market demand for Roman candles is given by Qd=1500-25P and marginal revenue curve by MR=20-Q/25 .Calculate the perfectly conpetitive and monopoly outputs and prices .


1
Expert's answer
2020-08-11T08:39:28-0400

The perfectly conpetitive output and price are:

Pd = MC,

P = 40 - 0.04Q,

"40 - 0.04Q = 12,"

Q = 700 units,

"P = 40 - 0.04*700 = 28."

The monopoly output and price are:

MC = MR,

"MR = TR'(Q) = 40 - 0.08Q,"

"40 - 0.08Q = 16,"

Q = 300 units,

"P = 40 - 0.04*300 = 28."